OpenStudy (anonymous):

solve for theta sqrt3csc(theta)-2cot(theta)=0

OpenStudy (anonymous):

this is a tough one any ideas?

OpenStudy (anonymous):

i cheated and looked up the answer doesn't look like anything that will be easy to find http://www.wolframalpha.com/input/?i=sqrt%283%29csc%28theta%29-2cot%28theta%29%3D0

OpenStudy (anonymous):

i turned everything into terms of sine and cos and i just stared at it for a bit :/

OpenStudy (anonymous):

haha that works too, thanks (:

OpenStudy (anonymous):

well doesn't really explain much it looks like it is cooked to be nice what with the \(\sqrt{3}\) and the \(2\) like perhaps it is supposed to be the result of some double angle formula, but then when you see the answer....

OpenStudy (anonymous):

\[ \begin{array}{rcr} \sqrt{3}\csc(\theta)-2\cot(\theta) & = & 0 \\ \frac{\sqrt{3}}{\sin(\theta)}-\frac{2\cos(\theta)}{\sin(\theta)} & = & 0 \\ \sqrt{3}-2\cos(\theta) & = & 0 \\ 2\cos(\theta) & = & \sqrt{3} \\ \cos(\theta) & = & \frac{\sqrt{3}}{2} \\ \theta & = & \frac{\pi}{6} \end{array} \] We assumed that \(\sin(\theta) \neq 0 \) as it would be undefined if that where true. Also since \(\cos(\theta) \) is periodic, there are many solutions, I just gave the most obviou s one.

OpenStudy (anonymous):

i guess this is what is asked for although apparently there are lots of other solutions

OpenStudy (anonymous):

Consider\[\cos(\theta+2\pi n) = \cos(\theta) \quad n \in \mathbb{Z}\]And\[ \cos(-\theta) = \cos(\theta) \]That is what I mean when I say there are many answers.